OpenStudy (anonymous):

Shawn works 40 hours a week, if he made 27,040 last year, how much was he paid per hour

OpenStudy (anonymous):

you divide how much he made by 40 then divide it again by 24\times 7

OpenStudy (geneticrockhopper247):

27,040/52=520. =Amount made per week. 520/40=13=amount made per hour

OpenStudy (anonymous):

yes i geuss